# AIR-HLoc: Adaptive Retrieved Images Selection for Efficient Visual Localisation

### v2

@@ IV-D Results
IV-D1 Cambridge Landmarks
TABLE I: Comparisons on Cambridge Landmarks dataset. We report the median translation and rotation errors (cm/) of different methods. We report the results of HLoc and AIR-HLoc using $k=10$.
@@ IV-D1 Cambridge Landmarks
IV-D3 Aachen Day-Night-v1.1
For both HLoc and AIR-HLoc, the three accuracy levels converge when $k\geq 20$. Subfigures 4 (a)-(b) and Figure 5 (a) shows that AIR-HLoc achieves almost the same accuracy of HLoc with up to 11% fewer retrieved images. We find that AIR-HLoc (NetVLAD) has slightly better accuracy than HLoc (NetVLAD) in (0.25m, $2^{\circ}$) when $k=20$, and AIR-HLoc (EigenPlaces) have slightly better accuracy than HLoc (EigenPlaces) in (0.25m, $2^{\circ}$) when $k=30$ in day-time test sequences. This shows that our AIR-HLoc can improve accuracy by reducing the noisiness introduced by high $k$ for easy queries.
@@
Figure 4: Percentage (%) of test frames high (0.25m, $2^{\circ}$), medium (0.5m, $5^{\circ}$), and low (5m, $10^{\circ}$) accuracy [41] (higher is better) for HLoc and AIR-HLoc against $k$. The average retrieved ratio ($k^{*}/k$) for all queries is shown in Figure 5 (a).
@@ IV-E Analysis
Figure 6: (a)- (b) Plots of mean rotation and translation errors of HLoc for easy, medium and hard queries of described in Section III-C against $k$ in Cambridge Landmarks. (c)- (d) The average RAYR, $\zeta=(\zeta_{T}(k),\zeta_{R}(k))$ of all query frames against $k$ from 1 to 30 in Cambridge Landmarks.
@@ IV-F System Efficiency and Application
We first conduct the runtime analysis using an NVIDIA GeForce GTX 4090 GPU on a desktop. To illustrate the bottleneck introduced by feature matching and its impact on mobile robot localisation tasks, we also evaluate the runtime on an embedded NVIDIA Jetson Orin platform with $64$ GB RAM. GTX 4090: We assess the runtime of the entire HLoc pipeline. Since global and local feature extraction for the reference image database can be performed offline, we focus on the runtime for query image feature extraction (runtime1), image retrieval (runtime2), feature matching (runtime3), and pose estimation (runtime4). Runtime1 and runtime4 are nearly constant per query, while runtime2 depends on the size of the image database. For King’s College, runtime1 + runtime2 + runtime4 is approximately 40 ms per query. We observe that runtime3 increases linearly with $k$, as shown in Figure 5 (b). Runtime3 is longer for Cambridge and Aachen-V1.1 compared to 7Scenes due to the higher resolution of images in these datasets [^2] We resize Aachen Day-Night-v1.1 images to a maximum size of 1600 pixels and Cambridge landmarks images to 1024 pixels. For 7Scenes, we retain the original resolution of 640x480 pixels.. The average runtime3 per image pair is 17 ms for 7Scenes, 40 ms for Cambridge, and 107 ms for Aachen-V1.1. Runtime3 dominates the pipeline’s execution time when $k\geq 10$, while most datasets require at least $k=10$ for optimal performance. At $k=10$, AIR-HLoc reduces the average runtime3 across all queries by 44 ms, 120 ms, and 110 ms for 7Scenes, Cambridge, and Aachen-V1.1, respectively. Jetson Orin: The average runtime3 per image pair is 510 ms for 7Scenes, 811 ms for Cambridge, and 1900 ms for Aachen-V1.1. At $k=10$, AIR-HLoc reduces the average runtime3 across all queries by 1326 ms, 2433 ms, and 2090 ms for 7Scenes, Cambridge, and Aachen-V1.1, respectively. It shows that our AIR-HLoc significantly reduces the computational time on a mobile robot platform.
@@ V CONCLUSIONS
The IR module in HLoc identifies the top-$k$ similar images from a database of reference images for a query image. Increasing $k$ enhances localisation robustness for challenging queries but also leads to increased computational cost for feature matching. Feature matching is the bottleneck of the runtime, increasing linearly with $k$. This paper introduces AIR-HLoc, a method that optimizes processing time by adaptively retrieve different number of images based on the similarity between query images and the reference image database. Extensive experiments conducted on three datasets demonstrate the efficacy of our algorithm, achieving a reduction by up to 30% in matching cost while maintaining SOTA accuracy. This paper provides novel insights into the relationship between global and local descriptors, offering a deeper understanding that can inform more effective $k$ selection and algorithm design for practical applications.

### v3

@@ IV-D Results
TABLE I: Comparisons on Cambridge Landmarks dataset. Median translation and rotation errors (cm/). We report the results of HLoc and AIR-HLoc using $k=10$.
@@ IV-D1 Cambridge Landmarks
Figure 4: Percentage (%) of test frames high (0.25m, $2^{\circ}$), medium (0.5m, $5^{\circ}$), and low (5m, $10^{\circ}$) accuracy [41] (higher is better) for HLoc and AIR-HLoc against $k$. The average retrieved ratio ($k^{*}/k$) is shown in Figure 5 (a).
IV-D3 Aachen Day-Night-v1.1
For both HLoc and AIR-HLoc, the three accuracy levels converge when $k\geq 20$. Subfigures 4 (a)-(b) and Figure 5 (a) show that AIR-HLoc achieves almost the same accuracy as HLoc with up to 11% fewer retrieved images. We find that AIR-HLoc achieves accuracy comparable to that of HLoc. AIR-HLoc (EP) achieves slightly better accuracy than HLoc (EP) in the (0.25m, $2^{\circ}$) range when $k=30$ in daytime test sequences. This demonstrates that AIR-HLoc can improve accuracy by reducing the noise introduced by a high $k$ value.
@@ IV-E Analysis
Figure 6: (a)- (b) Plots of mean rotation and translation errors of HLoc for easy, medium and hard queries of described in Section III-C against $k$ in Cambridge Landmarks. (c)- (d) The average MLIP, $\zeta=(\zeta_{T}(k),\zeta_{R}(k))$ of all query frames against $k$ from 1 to 30 in Cambridge Landmarks.
@@ IV-F System Efficiency and Application
We first conduct the runtime analysis using an NVIDIA GeForce RTX 4090 GPU on a desktop. To illustrate the bottleneck introduced by feature matching and its impact on mobile robot localisation tasks, we also evaluate the runtime on an embedded NVIDIA Jetson Orin platform with $64$ GB RAM. 4090: We assess the runtime of the entire HLoc pipeline. Since global and local feature extraction for the reference image database can be performed offline, we focus on the runtime for query image feature extraction (runtime1), image retrieval (runtime2), feature matching (runtime3), and pose estimation (runtime4). Runtime1 and runtime4 are nearly constant per query, while runtime2 depends on the size of the image database. For King’s College, runtime1 + runtime2 + runtime4 is approximately 40 ms per query. We observe that runtime3 increases linearly with $k$, as shown in Figure 5 (b). Runtime3 is longer for Cambridge and Aachen-V1.1 compared to 7Scenes due to the higher resolution of images in these datasets [^2] We resize Aachen Day-Night-v1.1 images to a maximum size of 1600 pixels and Cambridge landmarks images to 1024 pixels. For 7Scenes, we retain the original resolution of 640x480 pixels.. The average runtime3 per image pair is 17 ms for 7Scenes, 40 ms for Cambridge, and 107 ms for Aachen-V1.1. Runtime3 dominates the pipeline’s execution time when $k\geq 10$, while most datasets require at least $k=10$ for optimal performance. At $k=10$, AIR-HLoc reduces the average runtime3 across all queries by 44 ms, 120 ms, and 110 ms for 7Scenes, Cambridge, and Aachen-V1.1, respectively. Jetson Orin: The average runtime3 per image pair is 510 ms for 7Scenes, 811 ms for Cambridge, and 1900 ms for Aachen-V1.1. At $k=10$, AIR-HLoc reduces the average runtime3 across all queries by 1326 ms, 2433 ms, and 2090 ms for 7Scenes, Cambridge, and Aachen-V1.1, respectively. It shows that our AIR-HLoc significantly reduces the computational time on a mobile robot platform.
@@ V CONCLUSIONS
The IR module in HLoc retrieves the top-$k$ similar reference images for a given query. A larger $k$ improves localisation robustness for challenging queries but increases the computational cost, as feature matching— the main runtime bottleneck—scales linearly with $k$. This paper introduces AIR-HLoc, a method that optimizes processing time by adaptively retrieving a different number of images based on the similarity between query images and the reference image database. Extensive experiments conducted on three datasets demonstrate the efficacy of our algorithm, achieving up to a 30% reduction in matching cost while maintaining state-of-the-art (SOTA) accuracy. This paper provides novel insights into the relationship between global and local descriptors, offering a deeper understanding that can inform more effective $k$ selection and algorithm design for practical applications.